A certain two digit number has a value that is three times the sum of it's digits. The units digit is one more than three times the tens digit. Find the number. Please help me solve this!!
Answer by ankor@dixie-net.com(22740)

let a = the 10's digit
let b = the units
then
10a + b = "a certain two digit number
:
"A certain two digit number has a value that is three times the sum of it's digits."
10a + b = 3(a + b)
10a + b = 3a + 3b
10a - 3a = 3b - b
7a = 2b
:
" The units digit is one more than three times the tens digit."
b = 3a + 1
:
Substitute for b in the 1st equation
7a = 2(3a+1)
7a = 6a + 2
7a - 6a = 2
a = 2
;
I'll let you find b, check your solution in the 1st statement